OpenStudy (anonymous):

To solve the equation 8w + 4 = -12, you would first divide each side of the equation by 8 and then subtract 4 from each side. True False

OpenStudy (anonymous):

the instruction would give us \[w - \frac{ 7 }{ 2 } =-\frac{ 11 }{ 2 }\] if there is no other instructions like adding both side by \[\frac{ 7 }{ 2 }\] then its a false